# LocCa: Visual Pretraining with Location-aware Captioners

### v1

@@ 4.2.2 Referring Expression Segmentation
For referring expression segmentation, we extend the REC task by adding a suffix “Mask:” to the text, followed by the indexes of segmentation tokens. The segmentation tokens specify the precise shape of the segmentation mask within the bounding box that is dentified during the REC task. This process leverages a pretrained VQ-VAE [56] to convert the semantic masks to tokens, please refer to Appendix 0.A.2 for more details.
LocCa is adapted to RES under the same settings as in REC, using the “clean” combined training sets of RefCOCO, RefCOCO+, and RefCOCOg (c.f. Appendix 0.C). Specifically, we compare different frozen encoders and train a decoder from scratch. As shown in Table 2, LocCa’s vision encoder outperforms other encoders substantially, thereby providing further validation of its location-wise sensitivity. Moreover, we employ the full encoder-decoder LocCa model and fine-tune it for RES. Notably, LocCa achieves competitive results even compared to the state-of-the-art PaLI-3 model, albeit with considerably fewer parameters (0.6B vs. 5B). See Appendix 0.A.2 for details.
@@ 4.2.3 Holistic Image Understanding
It is also interesting to investigate the full potential of the LocCa model by fine-tuning it on holistic tasks with a small $3\times 10^{-6}$ learning rate without weight decay. To this end, we choose the widely used COCO image captioning task as an example. LocCa achieves competitive results of 138.0 CIDEr score with a standard 2242px. When increasing the transfer resolution to 6402px, the LocCaL model achieves 140.3 CIDEr score without CIDEr metric optimization.
@@ 4.2.4 Vision-Language Models with LocCa
We select a wide range of tasks to assess the models’ proficiency in understanding various visual concepts, including: image scene (COCO Caption), objects (VQAv2 and TallyQA [1]), visually-situated text (TextVQA [68], ST-VQA [6]) and general knowledge (OKVQA [51]). Remarkably, as shown in Table 4, LocCa consistently surpasses the Cap and CapPa vision encoders by a significant margin across all these tasks. Note that SigLIPL in Table 4 uses image patch size 16 while all the other models use patch size 14. SigLIPL is marked grey because it’s not directly comparable. The LocCaG model consistently outperforms SigLIPG across all the tasks. The underlying reason for this performance improvement is attributed to the pretraining strategy where LocCa is explicitly instructed with object locations. With the knowledge of object and textual regions, the model is better positioned to understand more challenging and subtle concepts like counting, visual relations, and word characters.
@@ 4.2.5 Object Detection
Both [59, 9] pretrain their encoder-decoder model in Objects365 [67] and we found this to be critical. In particular the task output format is different from the ones LocCa used during pretraining and the size of COCO is too small to learn it without overfitting. To measure the performance of the encoders, we opt to initialize the decoder from an Objects365 pretrained model.
[…]
We present the comparative results in Figure 2, where LocCa significantly outperforms all baselines in both mAP and AR, at both stages — before and after the application of reinforcement tuning. This performance aligns with our expectations regarding LocCa’s object-centric comprehension, attributed to the integration of additional location-aware pretraining.
@@ 4.2.6 Zero-shot Transfer
Following locked-image tuning [99], we freeze the pre-trained vision encoder and train a text encoder contrastively to perform zero-shot classification and zero-shot retrieval on downstream tasks. With an L/14 architecture and 3B examples LiT-tune duration, LocCaL achieves 77.1% 0-shot accuracy on ImageNet, which is competitive to the same size CapPaL [72] model’s 76.4% accuracy. On COCO retrieval tasks, LocCaL achieves 46.6% and 64.6% on text-to-image and image-to-text retrieval tasks respectively. It outperforms CapPaL’s 43.9% and 60.3% COCO retrieval results by a large margin.
Table 5: Fine-tuning vision backbones with a linear head [69] for semantic segmentation on ADE20k
@@ 4.2.7 Semantic Segmentation
We investigate the dense feature learning capabilities of LocCa by evaluating it for semantic segmentation on ADE20k [102] along with Cap and CapPa. To this end, we use the Segmenter framework [69] which attaches a linear layer to every patch embedding to predict the semantic label of that patch, and obtains the high-resolution semantic map by upsampling the low-resolution map. The results in Table 5 show that LocCa outperforms Cap and CapPa by about 2 mIoU points, and the Seg ViT-L baseline which is based on an ImageNet-21k-pretrained ViT-L by about 1 point.
@@ 4.2.8 Video Understanding
We follow the PaLI-3 [11] structural setup to evaluate LocCa and CapPa on video understanding tasks. Specifically, we use the image encoder to process each frame separately and concatenate all resulting tokens. A LiT-Decoder [3] is then tuned on a mixture of six video understanding tasks. Despite no video has been seen during pretraining, LocCa showcases the video understanding capabilities and outperforms CapPa on most of the datasets. In particular, for the video captioning task on VATEX [79] dataset, LocCa achieves a CIDER score of 65.0 vs 64.0 of CapPa. For the video QA task on MSVD [84] dataset, LocCa obtains an accuracy of 50.9 vs 50.0 of CapPa. Please refer to the Appendix 0.A.6 for more detailed results.
@@ 4.3 Qualitative Results
LocCa already showcases the superior performance of its visual encoder in object detection when equipped with a pretrained box decoder. In this section, we will present the raw LocCa model’s zero-shot capability to detect multiple objects using its own decoder, despite only a single object per example being observed during pretraining. To achieve this, we employ an appropriate sampling strategy when decoding the output from LocCa. This involves using the prompt of a single task prefix “GCap:” to instruct LocCa to predict the RoIs along with their labels. As shown in Figure 4, we observe that depending on the setting we get lower number of bounding boxes with correct class names, or higher number of boxes with class names which start to contain noise. We provide more visualizations and discussions in Appendix 0.B. Nevertheless, pre-trained LocCa model shows powerful capability after a short finetuning on a clean downstream dataset as shown in Table 1. Finding a decoding strategy which can output high number of boxes and quality labels at the same time is an open problem. We leave the decoding strategy as well as instruction tuning of a LocCa model on multiple clean datasets for future work.
Table 6: Ablation study on applying loss on Ref. and Gcap. tasks during training.
REF GCAP INet COCO VQAv2 GQA RefCOCO RefCOCO+
@@
Figure 4: Visualization of LocCaL’s zero-shot predictions on Visual Genome[37]: (a) the top row was sampled with temperature 0.5 and top-k (k=40) (b) the bottom row was sampled with temperature 1 (and without top-k). In both cases non-maximum suppression was applied on the decoded 30 samples. We can see that adding more noise increases the variety of decoded boxes and degrades the quality of labels.
@@ 4.4.1 Pretraining and transfer resolutions
We present results with different pretrain resolution and transfer resolution combinations in Figure 3 (a). To get a LocCa model with higher 384 pretrain resolution, we finetune the 224 resolution model using the target 384 resolution for 900M examples seen on the same dataset. In Figure 3 (a), we transfer both the 224 resolution model and the 384 resolution model to RefCOCOs using different transfer resolutions (224, 384, 640) as marked on the x-axis. We find that the higher pretrain model achieves slightly higher or competitive results compared to the 224 resolution models. We opt for the 224 resolution pretrain models by default in this paper for simplicity. More results with high pretrain resolution could be found in the Appendix 0.A.7.
@@ 4.4.3 Selection of Pretraining tasks
To evaluate the importance of the selected tasks for pretraining LocCa, an ablation study was conducted focusing on the effects of including REF and GCAP tasks. Specifically, the study explored the impact on LocCa by removing these tasks individually and collectively, with the model defaulting to the CapPa baseline when both are excluded. All these models are pretrained on the WebLI split for 900M examples with the resolution of 2242px, and subsequently evaluated on the holistic tasks using a LiT-Decoder setup and on RefCOCOs by fine-tuning the whole model.
@@
As shown in Table 6, incorporating any location-aware task into the pretraining of LocCa yields significant performance improvements, particularly evident in the RefCOCO results. Interestingly, incorporating solely the GCAP task leads to a marked improvement in RefCOCO performance compared to the CapPa baseline, despite GCAP not being directly aligned with the REF task. This highlights the importance of introducing object concepts for enhancing regional-level understanding, which is beneficial for fine-grained visual comprehension and applicable to other object-sensitive tasks. Furthermore, the combined inclusion of both REF and GCAP tasks yields even better results, demonstrating their complementary nature in improving LocCa’s performance.
5 Conclusion and Future works
@@
LocCa introduces a powerful approach to visual pretraining, combining simple, effective location-aware techniques with the power of auto-regressive decoders. This model excels in understanding both the overall scene and specific spatial details, setting new performance standards on location-aware tasks while preserving the capability on holistic image understanding. LocCa simplifies the process of blending location information with visual data, offering significant improvements on tasks requiring detailed spatial awareness. Our contributions pave the way for advanced model capabilities in processing a broad spectrum of vision-language tasks, demonstrating LocCa’s versatility and effectiveness.
@@
LocCa already demonstrates superior zero-shot detection capability in qualitative results; we leave zero-shot segmentation to future work. Building on the robust foundation of LocCa, a promising direction for future work involves enhancing its pixel-level precision through the incorporation of segmentation tasks during the pretraining phase. This extension aims to equip LocCa with a more nuanced understanding of images, enabling it to discern and interpret intricate details and textures with unparalleled accuracy, further broadening its applicability across diverse vision-language tasks.
5.0.1 Acknowledgements.
@@
We thank Matthias Minderer, Jeremiah Harmsen for their valuable feedback on the manuscript. We thank Debidatta Dwibedi, Xingyi Zhou for their feedback on dense captioning evaluators. We thank Paul Voigtlaender for discussions on the referring expression segmentation task. We thank the Google DeepMind team for providing a supportive research environment. We use the big_vision codebase [5, 4] for all experiments in this project.
Appendix 0.A Details on transfer results and ablations
@@
0.A.1 LiT-Decoder setup for Referring Expression Comprehension and holistic image understanding tasks
@@
0.A.2 Referring Expression Segmentation
@@
In Table 7, we showcase the results of employing the full encoder-decoder LocCa model and fine-tuning it for RES task. It achieves competitive results even compared to the state-of-the-art PaLI-3 model, with considerably fewer parameters.
@@
0.A.3 Vision-Language Models with LocCa
@@
0.A.4 Zero-shot transfer
@@ 0.A.5 Semantic segmentation on ADE20k
0.A.6 Experiments on video understanding
@@ C B C B A A A J
0.A.7 Impact of high resolution pretrain and tokenization strategies
@@
Figure 7: More visualizations of LocCaL’s zero-shot predictions on Visual Genome[37]: from left to right we increase noise when sampling. The top rows show all the boxes without non-maximum suppression and the bottom rows show captions for boxes passing the non-maximum suppression filter. We can see that adding more noise increases the variety and amount of decoded boxes and at the same time degrades the quality of captions.
Appendix 0.B More Visualization Results
We provide more visualizations of LocCaL’s zero-shot predictions on Visual Genome[37] in Fig. 7. We have two notable observations: (i) When comparing the top rows to the bottom rows, LocCa excels at identifying foreground regions, yet the box regions exhibit significant overlap without non-maximum suppression. This is because only one single object per example is observed during LocCa pretraining, leading to the selection of RoIs that are highly random. (ii) From left to right we increase noise during sampling, which increases the variety and quantity of decoded boxes, while simultaneously degrades the quality of captions. We hypothesize that box sampling demands higher noise levels to explore to explore more diverse RoIs, whereas text sampling requires lower noise levels to ensure the generated texts are highly semantic relevant.
Appendix 0.C Duplicate image ids for RefCOCOs

### v2

@@ 4.2.2 Referring Expression Segmentation
For referring expression segmentation, we extend the REC task by adding a suffix “Mask:” to the text, followed by the indexes of segmentation tokens. The segmentation tokens specify the precise shape of the segmentation mask within the bounding box that is identified during the REC task. This process leverages a pretrained VQ-VAE [86] to convert the semantic masks to tokens, please refer to Appendix A.3 for more details.
LocCa is adapted to RES under the same settings as in REC, using the “clean” combined training sets of RefCOCO, RefCOCO+, and RefCOCOg (c.f. Appendix C). Specifically, we compare different frozen encoders and train a decoder from scratch. As shown in Table 2, LocCa’s vision encoder outperforms other encoders substantially, thereby providing further validation of its location-wise sensitivity. Moreover, we employ the full encoder-decoder LocCa model and fine-tune it for RES. Notably, LocCa achieves competitive results even compared to the state-of-the-art PaLI-3 model, albeit with considerably fewer parameters (0.6B vs. 5B). See Appendix A.3 for details.
@@ 4.2.3 Holistic Image Understanding
It is also interesting to investigate the full potential of the LocCa model by fine-tuning it on holistic tasks with a small $3\times 10^{-6}$ learning rate without weight decay. To this end, we choose the widely used COCO image captioning task as an example. LocCa achieves competitive results of 138.0 CIDEr score with a standard 2242px. When increasing the transfer resolution to 6402px, the LocCaL model achieves 140.3 CIDEr score without CIDEr metric optimization [87].
@@ 4.2.4 Vision-Language Models with LocCa
We select a wide range of tasks to assess the models’ proficiency in understanding various visual concepts, including: image scene (COCO Caption), objects (VQAv2 and TallyQA [89]), visually-situated text (TextVQA [90], ST-VQA [91]) and general knowledge (OKVQA [92]). Remarkably, as shown in Table 4, LocCa consistently surpasses the Cap and CapPa vision encoders by a significant margin across all these tasks. Note that SigLIPL in Table 4 uses image patch size 16 while all the other models use patch size 14. SigLIPL is marked grey because it’s not directly comparable. The LocCaG model consistently outperforms SigLIPG across all the tasks. With the knowledge of object and textual regions, LocCa is better positioned to understand more challenging and subtle concepts like counting, visual relations, and word characters. Further fine-tuning details and the hyperparameters we used are provided in Appendix A.4.
@@ 4.2.5 Object Detection
Both [88, 50] pretrain their encoder-decoder model in Objects365 [93] and we found this to be critical. In particular the task output format is different from the ones LocCa used during pretraining and the size of COCO is too small to learn it without overfitting. To measure the performance of the encoders, we opt to initialize the decoder from an Objects365 pretrained model. More details of adapting LocCa for object detection please refer to Appendix A.5.
We present the comparative results in Figure 3, where LocCa significantly outperforms all image-text pretrained baselines in both mAP and AR, at both stages — before and after the application of reinforcement tuning. This performance aligns with our expectations regarding LocCa’s object-centric comprehension, attributed to the integration of additional location-aware pretraining.
@@ 4.2.6 Zero-shot Transfer
Following locked-image tuning [36], we freeze the pre-trained vision encoder and train a text encoder contrastively to perform zero-shot classification and zero-shot retrieval on downstream tasks. With an L/14 architecture and 3B examples LiT-tune duration, LocCaL achieves 77.1% 0-shot accuracy on ImageNet, which is competitive to the same size CapPaL [14] model’s 76.4% accuracy. On COCO retrieval tasks, LocCaL achieves 46.6% and 64.6% on text-to-image and image-to-text retrieval tasks respectively. It outperforms CapPaL’s 43.9% and 60.3% COCO retrieval results by a large margin. Please refer to Appendix A.6 for more details.
Table 5: Fine-tuning vision backbones with a linear head [94] for semantic seg. on ADE20k.
@@ 4.2.7 Semantic Segmentation
We investigate the dense feature learning capabilities of LocCa by evaluating it for semantic segmentation on ADE20k [95] along with Cap and CapPa. To this end, we use the Segmenter framework [94] which attaches a linear layer to every patch embedding to predict the semantic label of that patch, and obtains the high-resolution semantic map by upsampling the low-resolution map. The results in Table 5 show that LocCa outperforms Cap and CapPa by about 2 mIoU points, and the Seg ViT-L baseline which is based on an ImageNet-21k-pretrained ViT-L by about 1 point. The transfer details are provided in Appendix A.7.
@@ 4.2.8 Video Understanding
We follow PaLI-3 [31] to evaluate LocCa and CapPa on video understanding tasks. Specifically, we use the image encoder to process each frame separately and concatenate all resulting tokens. A LiT-Decoder [17] is then tuned on a mixture of six video understanding tasks. Despite no video has been seen during pretraining, LocCa showcases the video understanding capabilities and outperforms CapPa on most of the datasets. In particular, for the video captioning task on VATEX [96] dataset, LocCa achieves a CIDER score of 65.0 vs 64.0 of CapPa. For the video QA task on MSVD [97] dataset, LocCa obtains an accuracy of 50.9 vs 50.0 of CapPa. More details could be found from Appendix A.8.
@@ 4.3 Qualitative Results
In this section, we discuss the raw LocCa model’s zero-shot capability to detect multiple objects using its own decoder, despite only a single object per example being observed during pretraining. To achieve this, we employ beam search when decoding the output from LocCa. This involves using the prompt of a single task prefix “GCap:” to instruct LocCa to predict the RoIs along with their labels. As shown in Appendix Fig. 6, we observe that depending on the setting we get lower number of bounding boxes with correct class names, or higher number of boxes with class names which start to contain noise. We provide more discussions in Appendix B. Nevertheless, pre-trained LocCa model shows powerful capability after a short finetuning on a clean downstream dataset as shown in Table 1. Finding a decoding strategy which can output high number of boxes and quality labels at the same time is an open problem.
Table 6: Ablation study on applying loss on AREF and GCAP tasks during training.
AREF GCAP INet COCO VQAv2 GQA RefCOCO RefCOCO+
@@ 4.4.1 Pretraining and transfer resolutions
We present results with different pretrain resolution and transfer resolution combinations in Figure 3 (a). To get a LocCa model with higher 384 pretrain resolution, we finetune the 224 resolution model using the target 384 resolution for 900M examples seen on the same dataset. In Figure 3 (a), we transfer both the 224 resolution model and the 384 resolution model to RefCOCOs using different transfer resolutions (224, 384, 640) as marked on the x-axis. We find that the higher pretrain model achieves slightly higher or competitive results compared to the 224 resolution models. We opt for the 224 resolution pretrain models by default in this paper for simplicity. More results with high pretrain resolution could be found in the Appendix A.10.
@@ 4.4.3 Selection of pretraining tasks
To evaluate the importance of the selected tasks for pretraining LocCa, an ablation study was conducted focusing on the effects of AREF and GCAP tasks. Specifically, the study explored the impact on LocCa by removing these tasks individually and collectively, with the model defaulting to the CapPa baseline when both are excluded. All these models are pretrained on the WebLI split for 900M examples with the resolution of 2242px, and subsequently evaluated on the holistic tasks using a LiT-Decoder setup and on RefCOCOs by fine-tuning the whole model. As shown in Table 6, incorporating any location-aware task into the pretraining of LocCa yields significant performance improvements, particularly evident in the RefCOCO results. Interestingly, incorporating solely the GCAP task leads to a marked improvement in RefCOCO performance compared to the CapPa baseline, despite GCAP not being directly aligned with the AREF task. This highlights the importance of introducing object concepts for enhancing regional-level understanding, which is beneficial for fine-grained visual comprehension and applicable to other object-sensitive tasks. Furthermore, the combined inclusion of both AREF and GCAP tasks yields even better results, demonstrating their complementary nature in improving LocCa’s performance.
5 Conclusion
@@
LocCa introduces a novel visual pretraining paradigm, using natural language interface to construct the proxy location-aware pretrain tasks. This model excels in understanding both the overall scene and specific spatial details, setting new performance standards on location-aware tasks while preserving the capability on holistic image understanding. LocCa simplifies the process of blending location information with visual data, offering significant improvements on tasks requiring detailed spatial awareness. Our contributions pave the way for advanced model capabilities in processing a broad spectrum of vision-language tasks, demonstrating LocCa’s versatility and effectiveness.
@@
LocCa already demonstrates superior zero-shot detection capability in qualitative results. However, it lacks the capability for zero-shot segmentation due to the absence of pretraining on pixel-level annotations, and we leave this for future work. Building on the robust foundation of LocCa, a promising direction for future work involves enhancing its pixel-level precision through the incorporation of segmentation tasks during the pretraining phase. This extension aims to equip LocCa with a more nuanced understanding of images, enabling it to discern and interpret intricate details and textures with unparalleled accuracy, further broadening its applicability across diverse vision-language tasks.
Acknowledgements
@@
We thank Matthias Minderer and Jeremiah Harmsen for their valuable feedback on the manuscript, and Debidatta Dwibedi and Xingyi Zhou for their feedback on dense captioning evaluators. We also thank Paul Voigtlaender for discussions on the referring expression segmentation task. Finally, we thank the Google DeepMind team for providing a supportive research environment. We used the big_vision codebase [98, 99] for all experiments in this project.
Appendix A Details on transfer results and ablations
A.1 Transfer to downstream tasks
@@
A.2 LiT-Decoder setup for Referring Expression Comprehension and holistic image understanding tasks
@@
A.3 Referring Expression Segmentation
@@
In Table 7, we showcase the results of employing the full encoder-decoder LocCa model and fine-tuning it for RES task. It achieves competitive results even compared to the state-of-the-art PaLI-3 model, with considerably smaller model size.
@@
A.4 Vision-Language Models with LocCa
@@
A.5 Object Detection
[…]
A.6 Zero-shot transfer
@@ A.7 Semantic segmentation on ADE20k
A.8 Experiments on video understanding
@@ C B C B A A A J
A.8.1 Grounded captioning
For the grounded captioning task, we follow the standard setup that generates regional captions based on a given bounding box location [58, 108, 109]. We report the grounded captioning results on the Visual Genome dataset and compare them with state-of-the-art methods.
As shown in Table 10, LocCa (0.6B, without LLM), outperforms GPT4RoI [108] (7B, with LLM) on both METEOR and CIDEr scores. METEOR evaluates the precision, recall, and alignment of words between the generated and reference captions, while CIDEr assesses the similarity of n-grams between them. When compared with GLaMM [109] (7B, with LLM), LocCa performs better on METEOR but lags on CIDEr. This difference can be attributed to the relatively simple captions in Visual Genome, which typically consist of only a few words. LocCa uses a simpler decoder (0.3B params only) that closely matches the reference captions in terms of word choice and order, which aligns well with the dataset’s simple nature. GLaMM, on the other hand, uses a more complex LLM as its decoder, which likely generates more diverse captions that include n-grams that match the reference captions more effectively.
It is important to note that the pretrained LocCa encoder complements multimodal LLMs (i.e. as a better alternative option to the CLIP encoder), and we anticipate further performance gains on downstream tasks when combining both.
Table 10: Grounded captioning results on the Visual Genome dataset.
Model # Param mAP METEOR CIDEr
GRIT 1B 15.5 17.1 142
GPT4RoI 7B - 17.4 145.2
GLaMM 7B - 19.7 180.5
LocCaL 0.6B 34.5 20.7 157.0
A.9 Experiments on grounded captioning
A.10 Impact of high resolution pretraining and tokenization strategies
@@
Figure 6: Visualizations of LocCaL’s zero-shot predictions on Visual Genome[57]: from left to right we increase noise when sampling. The top rows show all the boxes without non-maximum suppression and the bottom rows show captions for boxes passing the non-maximum suppression filter. We can see that adding more noise increases the variety and amount of decoded boxes and at the same time degrades the quality of captions.
Appendix B More Visualization Results
We provide visualizations of LocCaL’s zero-shot predictions on Visual Genome[57] in Fig. 6. We have two notable observations: (i) When comparing the top rows to the bottom rows, LocCa excels at identifying foreground regions, yet the box regions exhibit significant overlap without non-maximum suppression. This is because only one single object per example is observed during LocCa pretraining, leading to the selection of RoIs that are highly random. (ii) From left to right we increase noise during sampling, which increases the variety and quantity of decoded boxes, while simultaneously degrades the quality of captions. We hypothesize that box sampling demands higher noise levels to explore to explore more diverse RoIs, whereas text sampling requires lower noise levels to ensure the generated texts are highly semantic relevant.
Appendix C Duplicate image ids for RefCOCOs